How is it when n goes to infinity,

T(n) = 4n*(1 - nlog4(3/4)) + nlog4(3)

becomes Big-Oh,

T(n) = O(n)

?
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

Let f(n)=4n*(1 - nlog4(3/4)) + nlog4(3)

g(n)=n

So, f(n)/g(n)=(4*n*(1 - nlog4(3/4))+nlog4(3))/n

=4*(1 - nlog4(3/4))+nlog4(3)-1

So,

for n tending inf

nlog4(3)-1 goes 0

4*(1 - nlog4(3/4)) goes inf

So,

f/g tends inf

Kindly revert for any queries

Thanks.